You can drive 70 mph on the freeway and 35 mph in town. If an 80 mile drive took you 2 hours, how long did you drive in town?
Found 2 solutions by solver91311, RAY100:
Answer by solver91311(24713)



Let be the distance driven in town. Let be the time driven in town. Then is the distance driven on the freeway, and is the time driven on the freeway.

The trip through town:



The trip on the freeway:



But since we can say:



Solve for

John


Answer by RAY100(1637)

Distance = Velocity * Time
.
Let f = freeway time, let t = town time
.
70 (f) + 35 (t) = 80
.
but f + t = 2,,,,,f = 2-t
.
70(2-t) +35 (t) = 80
.
140 -70t +35t = 80
.
-35 t = -60
.
t = 1.71 hrs
.
f = 2- 1.71 = .29 hrs
.
check
.
70 (.29) + 35(1.71) = 80,,,,ok